OpenStudy (anonymous):

12(x + 4) = 4(1 + 3x)

OpenStudy (ajprincess):

12(x + 4) = 4(1 + 3x) 12x+48=4+12x Can u solve for x nw?

OpenStudy (anonymous):

\[0 \neq 44\]

OpenStudy (anonymous):

isee no solution.ok thnkz